OpenStudy (goformit100):

Identify the slope and y-intercept of the inequality 9x + 3y > 6. slope = 3, y-intercept = 2 slope = -3, y-intercept = 2 slope = -3, y-intercept = 6 slope = negative 1 over 3, y-intercept = 2

OpenStudy (goformit100):

@kelliegirl33

OpenStudy (dls):

B

OpenStudy (dls):

9x + 3y > 6. Compare with y=mx+c for slope Put x=0 for y intercept

OpenStudy (goformit100):

Thank you sir

OpenStudy (anonymous):

9x + 3y > 6 3y > -9x + 6 y > -9/3x + 6/3 y > -3x + 2 slope = -3 and y intercept = 2